4x + 3y = -4
4x -y = 0

Answer by mananth(16949)

4x+3y=4 .............1
4x-y=0 .............2

multiply (1)by 1
Multiply (2) by 3
4x+3y=4
12x-3y =0

Add the two equations

16x=4

/16

x=0.25

plug value of x in (1)

4x+3y=4
1+3y=4
3y=4-1
3y=3
y=1